OpenStudy (anonymous):

I'm having problems with this integral. "Use polar coordinates to find the volume of the given solid - under the cone z=sqrt(x^2+y^2) and above the disk x^2+y^2<=4"

OpenStudy (anonymous):

We should use cylindrical coordinates for this integral. These are similar to polar coordinates with the addition of z coordinate. z goes from 0 to r, r goes from 0 to 2 and theta, the angle goes from 0 to 2*pi so the integral will be: \[\int\limits_{0}^{2\pi}\int\limits_{0}^{2}\int\limits_{0}^{r}r dz dr d \theta\] notice how you get an extra r in your integral when you go from regular coordinates to cylindrical coordinates.

OpenStudy (anonymous):

Well, since the problem says to use polar coordinates, I think we should try to do it that way. Some professors take off for using a different method.

OpenStudy (anonymous):

First thing to do is try to picture it. I'm working on a visual representation now.

OpenStudy (anonymous):

Let's check with trunsaid, they may not even be at triple integrals yet.

OpenStudy (anonymous):

I tried doing polar coordinates and I'm getting \[\int\limits_{0}^{2\pi}\int\limits_{r}^{2}rdrd \theta\]

OpenStudy (anonymous):

I don't see how you can determine the volume of three dimensional object with a two dimensional coordinate system.

OpenStudy (anonymous):

Yes, you can but there would be a function in integrand

OpenStudy (anonymous):

Refer to picture attached.

OpenStudy (anonymous):

Thomas, polar coordinates are 3 dimensional. There is a second angle.

OpenStudy (anonymous):

smoothmath, I thought that was spherical coordinates that has two angles?

OpenStudy (anonymous):

Yes, Phi and theta

OpenStudy (anonymous):

Well, spherical is an extension of polar.

OpenStudy (anonymous):

Okay, so theta is the angle with the x axis going around the z axis, okay? And Phi we'll call the angle with the x axis going around the y axis.

OpenStudy (anonymous):

With polar coordinates: \[\int\limits_{0}^{2\pi}\int\limits_{0}^{2}r*rdrd \theta\] this is the same as when you use cilindrical coordinates. There are now two r's in the inegrand one from the transformation and on for the height.

OpenStudy (anonymous):

Refer to picture for angles.

OpenStudy (anonymous):

Following, Trung?

OpenStudy (anonymous):

spherical coordinates are useless here, because you have cone and not a sphere of some kind.

OpenStudy (anonymous):

You're wrong, Thomas, and if you follow along, you'll see that.

OpenStudy (anonymous):

They actually make it really simple.

OpenStudy (anonymous):

@smoothman go for it

OpenStudy (anonymous):

theta goes from 0 to 2pi, obviously.

OpenStudy (anonymous):

Phi goes from 0 to arctan(2/4)

OpenStudy (anonymous):

And then r depends on phi. So it goes from 0 to 4/cos(phi)

OpenStudy (anonymous):

See it?

OpenStudy (anonymous):

\[\int\limits \int\limits \left(\sqrt{x^2+y^2}-\left(x^2+y^2-4\right)\right)dydx\] \[\int _0^{2\pi }\int _0^2\left(r-\left(r^2-4\right)\right)rdrd\theta \]

OpenStudy (anonymous):

I have a feeling trung is just gone, but hey.

OpenStudy (anonymous):

@smooth math, yea i understand how you did that

OpenStudy (anonymous):

@smoothman no, you might be correct, but it sure isn't making it easier.

OpenStudy (anonymous):

Well do you see the angles I drew? Theta goes from 0 to 2pi, right?

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

@imranmeah91, subtracting the two equations just took care of the change in R distance as you're integrating?

OpenStudy (anonymous):

Next we want to see the limits for phi. It starts at 0. Where does it end? The cone forms a right triangle with adjacent side 4 and opposite side length 2.

OpenStudy (anonymous):

So the angle for that is arctan(2/4). So phi goes from 0 to arctan(2/4)

OpenStudy (anonymous):

That's what we need to find volume, I believe

OpenStudy (anonymous):

This thread is getting messy, isn't the adjacent side 2?

OpenStudy (anonymous):

Yeah, to introduce spherical is to further complicate the problem and you are not following the teacher's instructions, which is to use polar coordinates. (Even if you say spherical is just an extension...) I think trunsaid should just follow imranmeah's example.

OpenStudy (anonymous):

Ah yes. Definitely. My mistake.

OpenStudy (anonymous):

thank you all for the help though. all of you made this whole concept of multiple integrals easier to understand

OpenStudy (anonymous):

Hmmm. You're right. You can do it in pure polar if you work with a third variable, z.

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^{2}r*rdrd \theta\] I still think this is correct

OpenStudy (anonymous):

z from 0 to 2, then theta from 0 to 2pi, then r from z to 2.

OpenStudy (anonymous):

Magic.

OpenStudy (anonymous):

My bad, guys. I was complicating things. I'm just too stubborn.

OpenStudy (anonymous):

honestly though using cylindrical coordinates made this much easier for me but I'm pretty sure my professor wanted me to use polar

OpenStudy (anonymous):

This is the front view

OpenStudy (anonymous):

Last picture, I promise. This should clarify things.

OpenStudy (anonymous):

I don't think yours quite does it, Imran. r isn't purely from 0 to 2.

OpenStudy (anonymous):

I don't see how you get that integrand imran.

OpenStudy (anonymous):

It's a three dimensional region, so you need a third variable, z. r depends on z.

OpenStudy (anonymous):

Thomas, does my approach make sense to you?

OpenStudy (anonymous):

z=r, so why not make the integrand r*r?

OpenStudy (anonymous):

@smoothman, sort of, but as I've said, z=r, so just put r in the integrand

OpenStudy (anonymous):

Thomas9 is right the height is \[\sqrt{x^2+y^2}\]=r add up the heights at all radius and angle \[\int _0^{2\pi }\int _0^2r\text{ }rdrd\theta \]

OpenStudy (anonymous):

I don't see why that wouldn't work.

OpenStudy (anonymous):

me neither

OpenStudy (anonymous):

doesn't Thomas9's way only make sense if the radius is constant?

OpenStudy (anonymous):

No, I think it works fine because you integrate for r from 0 to 2. You aren't treating it as a constant.

OpenStudy (anonymous):

My way I think is a little more intuitive, but I think his is right too.

OpenStudy (anonymous):

And only a double integral while mine is a triple, so that's cool.

OpenStudy (anonymous):

Indeed, r is not constant.

OpenStudy (anonymous):

I still think the spherical triple integral is a particularly elegant solution. =D